OpenStudy (anonymous):

] The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours. Age Ticket Price Under three years free Three to twelve years $32 Over twelve years $44 On a certain day, the company took 82 people on whale watching trips. There were 54 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

can u tell me how many people are there aged over 12

OpenStudy (anonymous):

it doesnt say obviously

OpenStudy (anonymous):

@kristopher1014

OpenStudy (anonymous):

the company took 82 people out of which 54 children aged twelve and under

OpenStudy (anonymous):

now can u?

OpenStudy (anonymous):

28?

OpenStudy (anonymous):

yep....

OpenStudy (anonymous):

Now, 28*44+ (54-x) 32 = C

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@kristopher1014

OpenStudy (anonymous):

how do i solve 54-x?

OpenStudy (anonymous):

(54-x)*32 = 54*32 - 32x

OpenStudy (anonymous):

so its 28*44+54*32 - 32x=c ?